# competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; Rational-Function Approximation

### v1

@@ Abstract
This paper explores how competing interactions in the intermolecular potential of fluids affect their structural transitions. The study employs a versatile potential model with a hard core followed by two constant steps, representing wells or shoulders, analyzed in both one-dimensional (1D) and three-dimensional (3D) systems. Comparing these dimensionalities highlights the effect of confinement on structural transitions. Exact results are derived for 1D systems, while the Rational-Function Approximation is used for unconfined 3D fluids. Both scenarios confirm that when the steps are repulsive, the wavelength of the oscillatory decay of the total correlation function evolves with temperature either continuously or discontinuously. In the latter case, a discontinuous oscillation crossover line emerges in the temperature–density plane. For an attractive first step and a repulsive second step, a Fisher–Widom line appears. Although the 1D and 3D results share common features, dimensionality introduces differences: these behaviors occur in distinct temperature ranges, require deeper wells, or become attenuated in 3D. Certain features observed in 1D may vanish in 3D. We conclude that fluids with competing interactions exhibit a rich and intricate pattern of structural transitions, demonstrating the significant influence of dimensionality and interaction features.
@@
competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; Rational-Function Approximation
@@ 1 Introduction
It is well known that both statistical mechanics and thermodynamics aim at explaining the same phenomena concerning, among other issues, energy, work, and heat exchange in the different systems. While the first approach involves a purely microscopic approximation, the second one is macroscopic in nature. Nevertheless, one of the major purposes of statistical physics is the interpretation and prediction of the macroscopic properties of a system in terms of the interactions between its particles. In the case of liquids, one attempts to understand why and under what circumstances certain phases are stable in well-defined intervals of density and temperature and also to try to relate the thermodynamic, structural, and dynamic properties of those phases with the form and size of the molecules that form the liquid and the nature of the intermolecular interactions McQuarrie (1976).
@@
For the description of a multibody system such as a liquid, it is often enough to consider simplified representations which are able to capture the essential elements of real interactions and lead to an adequate description of the observed phenomenology. Therefore, the great attention that has been paid during many decades to interaction potentials consisting of a hard core followed by one or many piecewise constant sections of different widths and heights (which include the square-well and the square-shoulder potentials), is not surprising Chapela et al. (1984, 1989); Benavides and Gil-Villegas (1999); Franzese et al. (2001); Malescio et al. (2002); Skibinsky et al. (2004); Malescio et al. (2005); Benavides et al. (2006); Cervantes et al. (2007); Guillén-Escamilla et al. (2007); Rżysko et al. (2008); de Oliveira et al. (2008); de Oliveira et al. (2009); Rżysko et al. (2010); Hlushak et al. (2011); Yuste et al. (2011); Bárcenas et al. (2011); Loredo-Osti and Castañeda-Priego (2012); Santos et al. (2012); Bárcenas et al. (2013); Santos et al. (2013); Kim and Kim (2013); Yuste et al. (2022); Perdomo-Pérez et al. (2022); Largo and Solana (2014). With this class of potentials, it has been possible to model and understand many phenomena, such as liquid-liquid transitions Skibinsky et al. (2004); Malescio et al. (2005); Cervantes et al. (2007), colloidal interactions Guillén-Escamilla et al. (2007), the density anomaly in water and supercooled liquids de Oliveira et al. (2008); de Oliveira et al. (2009), and the thermodynamic and transport properties of Lennard-Jones fluids Chapela et al. (1984, 1989). In particular, in the case of colloidal dispersions the interaction between a pair of macromolecules is modeled through an effective potential having a short-range attractive part and a long-range repulsive part Ruiz-Franco and Zaccarelli (2021); Kravtsiv et al. (2022); Carretas-Talamante et al. (2023). The competition between both parts of this potential leads to interesting phenomenology and induces changes in phase behavior and in the thermodynamic, structural, and transport properties of the system Perdomo-Pérez et al. (2022); Tan et al. (2024). Similarly, in the case of complex fluids, such competing interactions are associated with the aggregation or clustering of surfactants, macromolecules, and colloidal particles in solution, which in turn may produce self-assembly and microphase segregation Imperio and Reatto (2006); Archer and Wilding (2007); Bomont et al. (2012); Bollinger and Truskett (2016a, b); Hu and Charbonneau (2018); Malescio and Sciortino (2020); Bomont et al. (2020); Guillén-Escamilla et al. (2022); Munaò et al. (2022a, b); Costa et al. (2023); Bomont et al. (2024).
Although there is already a vast amount of work concerning the thermodynamic properties of fluids whose molecules interact through potentials in which the attractive and repulsive parts are in competition, the knowledge pertaining to the influence of such competitions on their structural properties is relatively scarce Kim and Kim (2013).
@@
The decay of the total correlation function, $h(r)=g(r)-1$, where $g(r)$ is the radial distribution function, serves as a crucial indicator of structural transitions in fluids. These transitions, whether oscillatory or monotonic, reflect changes in the spatial arrangement of particles, driven by the interplay between attractive and repulsive forces in the intermolecular potential. Understanding the decay of $h(r)$ has significant implications for phenomena such as crystallization, phase separation, self-assembly, and the mechanical properties of materials.
@@
All of this serves as a motivation for the present paper. In previous work, we have used the so-called Rational-Function Approximation (RFA) approach López de Haro et al. (2008); Santos (2016) to study various three-dimensional (3D) fluids whose intermolecular potentials consist of a hard-core followed by piecewise constant sections Yuste et al. (2011); Santos et al. (2012, 2013). This includes not only square-well and square-shoulder fluids but also systems where the intermolecular potential combines square shoulders and square wells.Yuste et al. (2022). We have also carried out studies of the asymptotic behavior of the direct and total correlation functions of binary hard-sphere fluid mixtures Pieprzyk et al. (2020, 2021a, 2021b), which, among other things, exhibit interesting phenomenology concerning structural transitions.
@@
In this paper, we aim to illustrate the effect of competing interactions on structural transitions in fluids. To this end, we consider a fluid of number density $\rho$ and absolute temperature $T$, where the intermolecular pair potential is given by:
@@
This potential includes a hard core of diameter $\sigma$ and two steps characterized by heights $\epsilon_{1}$ and $\epsilon_{2}$ and widths $(\lambda_{1}-1)\sigma$ and $(\lambda_{2}-\lambda_{1})\sigma$, respectively. The parameters $\lambda_{1}$ and $\lambda_{2}$ are constants satisfying $1<\lambda_{1}<\lambda_{2}$, where $\lambda_{2}\sigma$ denotes the total range of the potential. The sign of each $\epsilon_{j}$ ($j=1,2$) determines whether the corresponding step is a shoulder ($\epsilon_{j}>0$) or a well ($\epsilon_{j}<0$). This form of the potential is flexible enough to explore various competing interactions. In particular, when $\epsilon_{1}=\epsilon_{2}$ or $\epsilon_{2}=0$, the potential reduces to either the square-shoulder potential (for $\epsilon_{1}>0$) or the square-well potential (for $\epsilon_{1}<0$), making these cases particular limits of the general model.
@@
This work focuses on examining the qualitative changes in the structural behavior of the system as the potential transitions from the square-shoulder case to more complex potentials, where the second section is always a repulsive barrier ($\epsilon_{2}>0$).
@@
If both $\epsilon_{1}$ and $\epsilon_{2}$ are positive, the total correlation function $h(r)$ is expected to exhibit oscillatory decay. At very low temperatures, this decay has a wavelength on the order of the range of the repulsive barrier ($\lambda_{2}\sigma$). Conversely, at very high temperatures, the wavelength aligns with the hard-core diameter ($\sigma$). At a given density, the transition between these behaviors can occur either continuously or discontinuously. In the latter scenario, a discontinuous oscillation crossover (DOC) line would emerge, akin to the one observed in binary hard-sphere mixtures Pieprzyk et al. (2020, 2021a); Grodon et al. (2004, 2005); Statt et al. (2016); Royall et al. (2024).
@@
On the other hand, if $\epsilon_{1}<0$ and $\epsilon_{2}>0$, one might expect the presence of a Fisher–Widom (FW) line, which separates a region in the $T$ vs $\rho$ plane where the asymptotic decay of $h(r)$ is damped oscillatory from a region where the decay is purely exponential and monotonic. For a given $\epsilon_{2}>0$, a competition between a DOC line and an FW line could arise as $\epsilon_{1}$ transitions from positive to increasingly negative values.
@@
From this point onward, we adopt the hard-core diameter as the unit of length ($\sigma=1$), so all distances will be expressed in units of $\sigma$. The reduced density is then given by $\rho^{*}=\rho\sigma^{d}$, where $d$ is the dimensionality of the system. Since we assume $\epsilon_{2}>0$ throughout, we use $\epsilon_{2}$ as the unit of energy and define the reduced temperature as $T^{*}=k_{B}T/\epsilon_{2}$, with $k_{B}$ being the Boltzmann constant. However, when analyzing the impact of the second barrier on the FW line (in cases where $\epsilon_{1}<0$), we also introduce a second reduced temperature, $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$, to capture the relevant energy scale. The key dimensionless parameters characterizing the potential are thus $\lambda_{1}$, $\lambda_{2}$, and the ratio $\epsilon_{1}/\epsilon_{2}$.
@@
For reasons that will become apparent later, we restrict the value of $\lambda_{2}$ to be less than or equal to 2. For symmetry considerations, we will generally fix $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$, except in cases where $\epsilon_{1}=\epsilon_{2}$, where the effect of $\lambda_{2}$ on the DOC line will be specifically examined.
@@
Finally, we note that both one-dimensional (1D) and three-dimensional (3D) fluids interacting via the potential $\varphi(r)$, as defined in Equation (1), will be examined in the following analysis. This dual approach allows us to explore the impact of strong confinement on the structural transitions in fluids with competing interaction potentials. The results for the 1D system will be derived from the exact general solution, while for the unconfined 3D system, we will employ the RFA.
@@
The paper is organized as follows. In Section 2, we consider the 1D fluid, for which exact results can be derived. This is followed in Section 3 by the parallel analysis of the unconfined 3D fluid, where a brief but self-contained description of the RFA method is provided. Section 4 concludes the paper with a discussion of the results, including the differences in the structural behavior of 1D and 3D fluids modeled with the same interaction potential, along with some concluding remarks. Mathematical details are presented in two Appendices.
2 The 1D system. Exact results
@@ 2.1 Theoretical background
We begin by considering a system confined to a 1D geometry. In this case, we can take advantage of the fact that Equation (1) satisfies the conditions that, for 1D fluids, lead to exact results for the thermodynamic and structural properties Santos (2016), namely that $\lim_{r\to 0}\varphi(r)=\infty$, $\lim_{r\to\infty}\varphi(r)=0$, and that each particle interacts only with its two nearest neighbors when $\lambda_{2}\leq 2$.
As in previous works on 1D fluids Salsburg et al. (1953); Lebowitz and Zomick (1971); Percus (1976, 1982); Heying and Corti (2004); Santos (2007); Ben-Naim and Santos (2009); Fantoni and Santos (2017); Montero and Santos (2019); Maestre and Santos (2020); Montero et al. (2024), it is convenient to work with the Laplace transforms of both the radial distribution function $g(r)$ and the Boltzmann factor $e^{-\beta\varphi(r)}$ (where $\beta\equiv 1/k_{B}T$). These transforms are, respectively, defined as
@@
In fact, working in the isothermal-isobaric ensemble, one can express $G(s)$ in terms of $\Omega(s)$ as Santos (2016)
@@
where $p$ is the pressure and $\Omega^{\prime}(s)\equiv\partial_{s}\Omega(s)=-\int_{0}^{\infty}\mathrm{d}r\,e^{-rs}re^{-\beta\varphi(r)}$. Furthermore, the density of the fluid is also related to $\Omega(s)$ and reads
@@
In principle, the total correlation function $h(r)$ can be expressed in terms of the infinite set of poles $\{s_{n}\}$ of $G(s)$, which correspond to the nonzero roots of $\Omega(s+\beta p)=\Omega(\beta p)$. These poles have negative real parts and may be either real ($s_{n}=-\kappa_{n}$) or form complex-conjugate pairs ($s_{n}=-\zeta_{n}\pm\imath\omega_{n}$). For simplicity, we will use the term “pole” to refer collectively to both real values and complex-conjugate pairs. The locations of these poles depend on the thermodynamic state, with the pole whose real part is closest to zero governing the asymptotic behavior of the total correlation function.
@@
where $\delta_{n}$ is the argument of the associated residue $|A_{\zeta_{n}}|e^{\pm\imath\delta_{n}}$. The first term on the right-hand side of Equation (5) dominates over the second one if $\zeta_{1}<\zeta_{2}$; conversely, the second term dominates if $\zeta_{1}>\zeta_{2}$. Given a value of $\beta p$, there may exist a certain temperature at which the conditions $\zeta_{1}=\zeta_{2}$ and $\omega_{1}\neq\omega_{2}$ are satisfied. The set of such states plotted on the $T$ vs $\beta p$ plane (or equivalently on the $T$ vs $\rho$ plane) defines the DOC line. When this line is crossed, the wavelength of the damped oscillations in $h(r)$ undergoes a discontinuous shift from $2\pi/\omega_{1}$ to $2\pi/\omega_{2}$ (or vice versa).
@@
For a given value of $\beta p$, there may exist a specific temperature at which the conditions $\zeta=\kappa$ and $\omega\neq 0$ are satisfied. The collection of such states, when plotted on the $T$ vs $\beta p$ plane (or equivalently on the $T$ vs $\rho$ plane), defines the FW line. Upon crossing this line, the nature of the decay of the total correlation function $h(r)$ transitions between damped oscillatory and monotonic behavior (or vice versa).
@@
Regardless of the sign of $\epsilon_{j}$, the leading pole at a given density $\rho^{*}$ in the high-temperature limit ($\beta\to 0$) is given by $\zeta=\zeta_{\text{HR}}(\rho^{*})$ and $\omega=\omega_{\text{HR}}(\rho^{*})$, as shown in Appendix A.2, where the subscript HR refers to the hard-rod fluid. The HR oscillation frequency satisfies $\frac{1}{2}<\omega_{\text{HR}}(\rho^{*})/2\pi<1$, with the lower and upper bounds corresponding to $\rho^{*}\to 0$ and $\rho^{*}\to 1$, respectively. On the other hand, if $\epsilon_{j}>0$ and $\rho^{*}<\lambda_{2}^{-1}$, the leading pole in the low-temperature limit ($\beta\to\infty$) is given by $\zeta=\lambda_{2}^{-1}\zeta_{\text{HR}}(\rho^{*}\lambda_{2})$ and $\omega=\lambda_{2}^{-1}\omega_{\text{HR}}(\rho^{*}\lambda_{2})$ (see Appendix A.3.1). However, the low-temperature limit for $\lambda_{2}^{-1}<\rho^{*}<1$ is more intricate, as detailed in Appendix A.3.2.

### v2

# competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; rational function approximation

@@ Abstract
This paper explores how competing interactions in the intermolecular potential of fluids affect their structural transitions. This study employs a versatile potential model with a hard core followed by two constant steps, representing wells or shoulders, analyzed in both one-dimensional (1D) and three-dimensional (3D) systems. Comparing these dimensionalities highlights the effect of confinement on structural transitions. Exact results are derived for 1D systems, while the rational function approximation is used for unconfined 3D fluids. Both scenarios confirm that when the steps are repulsive, the wavelength of the oscillatory decay of the total correlation function evolves with temperature either continuously or discontinuously. In the latter case, a discontinuous oscillation crossover line emerges in the temperature–density plane. For an attractive first step and a repulsive second step, a Fisher–Widom line appears. Although the 1D and 3D results share common features, dimensionality introduces differences: these behaviors occur in distinct temperature ranges, require deeper wells, or become attenuated in 3D. Certain features observed in 1D may vanish in 3D. We conclude that fluids with competing interactions exhibit a rich and intricate pattern of structural transitions, demonstrating the significant influence of dimensionality and interaction features.
@@
competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; rational function approximation
@@ 1 Introduction
It is well known that both statistical mechanics and thermodynamics aim at explaining the same phenomena concerning, among other issues, energy, work, and heat exchange in different systems. While the first approach involves a purely microscopic approximation, the second one is macroscopic in nature. Nevertheless, one of the major purposes of statistical physics is the interpretation and prediction of the macroscopic properties of a system in terms of the interactions between its particles. In the case of liquids, one attempts to understand why and under what circumstances certain phases are stable in well-defined intervals of density and temperature and also to try to relate the thermodynamic, structural, and dynamic properties of those phases with the form and size of the molecules that form the liquid and the nature of the intermolecular interactions McQuarrie (1976).
@@
For the description of a multibody system such as a liquid, it is often enough to consider simplified representations which are able to capture the essential elements of real interactions and lead to an adequate description of the observed phenomenology. Therefore, the great attention that has been paid during many decades to interaction potentials consisting of a hard core followed by one or many piecewise constant sections of different widths and heights (which include the square-well and the square-shoulder potentials) is not surprising Chapela et al. (1984, 1989); Benavides and Gil-Villegas (1999); Franzese et al. (2001); Malescio et al. (2002); Skibinsky et al. (2004); Malescio et al. (2005); Benavides et al. (2006); Cervantes et al. (2007); Guillén-Escamilla et al. (2007); Rżysko et al. (2008); de Oliveira et al. (2008); de Oliveira et al. (2009); Rżysko et al. (2010); Hlushak et al. (2011); Yuste et al. (2011); Bárcenas et al. (2011); Loredo-Osti and Castañeda-Priego (2012); Santos et al. (2012); Bárcenas et al. (2013); Santos et al. (2013); Kim and Kim (2013); Yuste et al. (2022); Perdomo-Pérez et al. (2022); Largo and Solana (2014). With this class of potentials, it has been possible to model and understand many phenomena, such as liquid–liquid transitions Skibinsky et al. (2004); Malescio et al. (2005); Cervantes et al. (2007), colloidal interactions Guillén-Escamilla et al. (2007), the density anomaly in water and supercooled liquids de Oliveira et al. (2008); de Oliveira et al. (2009), and the thermodynamic and transport properties of Lennard–Jones fluids Chapela et al. (1984, 1989). In particular, in the case of colloidal dispersions, the interaction between a pair of macromolecules is modeled through an effective potential with a short-range attractive part and a long-range repulsive part Ruiz-Franco and Zaccarelli (2021); Kravtsiv et al. (2022); Carretas-Talamante et al. (2023). The competition between both parts of this potential leads to an interesting phenomenology and induces changes in phase behavior and in the thermodynamic, structural, and transport properties of the system Perdomo-Pérez et al. (2022); Tan et al. (2024). Similarly, in the case of complex fluids, such competing interactions are associated with the aggregation or clustering of surfactants, macromolecules, and colloidal particles in solution, which in turn may produce self-assembly and microphase segregation Imperio and Reatto (2006); Archer and Wilding (2007); Bomont et al. (2012); Bollinger and Truskett (2016a, b); Hu and Charbonneau (2018); Malescio and Sciortino (2020); Bomont et al. (2020); Guillén-Escamilla et al. (2022); Munaò et al. (2022a, b); Costa et al. (2023); Bomont et al. (2024).
There is an extensive body of research on the thermodynamic and structural properties of fluids whose molecules interact via competing attractive and repulsive forces. Particular attention has been given to systems described by variants of the short-range attraction and long-range repulsion (SALR) potential. These include models such as the two-Kac potential, the double Yukawa potential, the Lennard-Jones potential followed by a repulsive Yukawa tail, and the square-well potential followed by a repulsive ramp. For examples and further details, see Refs. Denton and Löwen (1997); Sear and Gelbart (1999); Pini et al. (2000); Mossa et al. (2004); Sciortino et al. (2004); Archer et al. (2007); Archer and Wilding (2007); Archer et al. (2008); Bomont and Costa (2012); Loredo-Osti and Castañeda-Priego (2012); Pȩkalski et al. (2013); Kim and Kim (2014); Godfrin et al. (2014); Sweatman et al. (2014); Cigala et al. (2015); Zhuang and Charbonneau (2016); Liu and Xi (2019); Bretonnet (2019); Charbonneau and Kundu (2020); Ruiz-Franco and Zaccarelli (2021); Valadez-Pérez et al. (2021); Costa et al. (2023).
In colloidal systems, the competition between short-range attraction and long-range repulsion leads to the emergence of intermediate-range-order structures, resulting in the formation of stable periodic microphases. This competition also disrupts the liquid–vapor phase transition, with the specific form of the SALR potential significantly influencing the morphology of the resulting structures Loredo-Osti and Castañeda-Priego (2012); Liu and Xi (2019). The intermediate-range order is closely linked to a peak in the static structure factor $S(k)$ (where $k$ is the wavenumber). Specifically, a divergence of $S(k)$ at $k=0$ indicates an instability associated with large-scale density fluctuations, while a divergence at a finite wavenumber signifies the presence of periodic microphases.
@@
Despite the extensive research on fluids whose molecules interact via competing attractive and repulsive forces, certain aspects of structural transitions in these systems remain unexplored. The decay of the total correlation function, $h(r)=g(r)-1$, where $g(r)$ is the radial distribution function, serves as a key indicator of such transitions. These transitions, characterized by oscillatory or monotonic decay, reflect changes in the spatial arrangement of particles arising from the delicate balance between attraction and repulsion in the intermolecular potential. A deeper understanding of the decay behavior of $h(r)$ is crucial for unraveling phenomena such as crystallization, phase separation, self-assembly, and the mechanical properties of complex materials.
@@
All this serves as a motivation for the present paper. In previous work, we have used the so-called rational function approximation (RFA) approach López de Haro et al. (2008); Santos (2016) to study various three-dimensional fluids whose intermolecular potentials consist of a hard core followed by piecewise constant sections Yuste et al. (2011); Santos et al. (2012, 2013). This includes not only square-well and square-shoulder fluids but also systems where the intermolecular potential combines square shoulders and square wells Yuste et al. (2022). We have also carried out studies of the asymptotic behavior of the direct and total correlation functions of binary hard-sphere fluid mixtures Pieprzyk et al. (2020, 2021), which, among other things, exhibit interesting phenomenology concerning structural transitions.
@@
In this paper, we aim to illustrate the effect of competing interactions on structural transitions in fluids. To this end, we consider a fluid of the number density $\rho$ and absolute temperature $T$, where the intermolecular pair potential is given by
@@
This potential includes a hard core of the diameter $\sigma$ and two steps characterized by the heights $\epsilon_{1}$ and $\epsilon_{2}$ and widths $(\lambda_{1}-1)\sigma$ and $(\lambda_{2}-\lambda_{1})\sigma$, respectively. The parameters $\lambda_{1}$ and $\lambda_{2}$ are constants satisfying $1<\lambda_{1}<\lambda_{2}$, where $\lambda_{2}\sigma$ denotes the total range of the potential. The sign of each $\epsilon_{j}$ ($j=1,2$) determines whether the corresponding step is a shoulder ($\epsilon_{j}>0$) or a well ($\epsilon_{j}<0$). This form of the potential is flexible enough to explore various competing interactions. In particular, when $\epsilon_{1}=\epsilon_{2}$ or $\epsilon_{2}=0$, the potential reduces to either the square-shoulder potential (for $\epsilon_{1}>0$) or the square-well potential (for $\epsilon_{1}<0$), making these cases particular limits of the general model. Studies on certain thermodynamic and structural properties of fluids whose molecules interact via a potential of the form given in Equation (1) have been reported in Refs. Denton and Löwen (1997); Loredo-Osti and Castañeda-Priego (2012); Kim and Kim (2013). However, it is important to note that, in our case, the range of the repulsive interaction is relatively short and cannot be accurately described as long-range.
@@
This work focuses on examining the qualitative changes in the structural behavior of a system as the potential transitions from the square-shoulder case to more complex potentials, where the second section is always a repulsive barrier ($\epsilon_{2}>0$).
@@
If both $\epsilon_{1}$ and $\epsilon_{2}$ are positive, the total correlation function $h(r)$ is expected to exhibit oscillatory decay. At very low temperatures, this decay has a wavelength in the order of the range of the repulsive barrier ($\lambda_{2}\sigma$). Conversely, at very high temperatures, the wavelength aligns with the hard core diameter ($\sigma$). At a given density, the transition between these behaviors can occur either continuously or discontinuously. In the latter scenario, a discontinuous oscillation crossover (DOC) line would emerge, akin to the one observed in binary hard-sphere mixtures Pieprzyk et al. (2020, 2021); Grodon et al. (2004, 2005); Statt et al. (2016); Royall et al. (2024).
@@
On the other hand, if $\epsilon_{1}<0$ and $\epsilon_{2}>0$, one might expect the presence of a Fisher–Widom (FW) line, which separates a region in the $T$ vs. $\rho$ plane where the asymptotic decay of $h(r)$ is damped in an oscillatory way from a region where the decay is purely exponential and monotonic. For a given $\epsilon_{2}>0$, a competition between a DOC line and an FW line could arise as $\epsilon_{1}$ transitions from positive to increasingly negative values.
@@
From this point onward, we adopt the hard core diameter as the unit of length ($\sigma=1$), so all distances will be expressed in units of $\sigma$. The reduced density is then given by $\rho^{*}=\rho\sigma^{d}$, where $d$ is the dimensionality of the system. Since we assume $\epsilon_{2}>0$ throughout, we use $\epsilon_{2}$ as the unit of energy and define the reduced temperature as $T^{*}=k_{B}T/\epsilon_{2}$, with $k_{B}$ being the Boltzmann constant. However, when analyzing the impact of the second barrier on the FW line (in cases where $\epsilon_{1}<0$), we also introduce a second reduced temperature, $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$, to capture the relevant energy scale. The key dimensionless parameters characterizing the potential are thus $\lambda_{1}$, $\lambda_{2}$, and the ratio $\epsilon_{1}/\epsilon_{2}$.
@@
For reasons that will become apparent later, we restrict the value of $\lambda_{2}$ to be less than or equal to 2. For symmetry considerations, we generally fix $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$, except in cases where $\epsilon_{1}=\epsilon_{2}$, where the effect of $\lambda_{2}$ on the DOC line is specifically examined.
@@
Finally, we note that both one-dimensional (1D) and three-dimensional (3D) fluids interacting via the potential $\varphi(r)$, as defined in Equation (1), will be examined in the following analysis. This dual approach allows us to explore the impact of strong confinement on structural transitions in fluids with competing interaction potentials. The results for the 1D system will be derived from the exact general solution, while for the unconfined 3D system, we will employ the RFA.
@@
The paper is organized as follows. In Section 2, we consider a 1D fluid. This is followed in Section 3 by the parallel analysis of an unconfined 3D fluid, where a brief but self-contained description of the RFA method is provided. Section 4 concludes the paper with a discussion of the results, including the differences in the structural behavior of 1D and 3D fluids modeled with the same interaction potential, along with some concluding remarks. Mathematical details are presented in Appendices A and B.
2 The 1D System: Exact Results
@@ 2.1 Theoretical Background
We begin by considering a system confined to a 1D geometry. In this case, we can take advantage of the fact that Equation (1) satisfies the conditions that, for 1D fluids, lead to exact results for thermodynamic and structural properties Santos (2016), namely that $\lim_{r\to 0}\varphi(r)=\infty$ and $\lim_{r\to\infty}\varphi(r)=0$ and that each particle interacts only with its two nearest neighbors when $\lambda_{2}\leq 2$.
As in previous works on 1D fluids Salsburg et al. (1953); Lebowitz and Zomick (1971); Percus (1976, 1982); Heying and Corti (2004); Montero and Santos (2019); Montero et al. (2024), it is convenient to work with the Laplace transforms of both the radial distribution function $g(r)$ and the Boltzmann factor $e^{-\beta\varphi(r)}$ (where $\beta\equiv 1/k_{B}T$). These transforms are, respectively, defined as
@@
In fact, working in the isothermal–isobaric ensemble, one can express $G(s)$ in terms of $\Omega(s)$ as Santos (2016)
@@
where $p$ is the pressure and $\Omega^{\prime}(s)\equiv\partial_{s}\Omega(s)=-\int_{0}^{\infty}\mathrm{d}r\,e^{-rs}re^{-\beta\varphi(r)}$. Furthermore, the density of the fluid is also related to $\Omega(s)$ and reads as
@@
In principle, the total correlation function $h(r)$ can be expressed in terms of the infinite set of poles $\{s_{n}\}$ of $G(s)$, which correspond to the nonzero roots of $\Omega(s+\beta p)=\Omega(\beta p)$. These poles have negative real parts and may be either real ($s_{n}=-\kappa_{n}$) or form complex–conjugate pairs ($s_{n}=-\zeta_{n}\pm\imath\omega_{n}$). For simplicity, we will use the term “pole” to refer collectively to both real values and complex–conjugate pairs. The locations of these poles depend on the thermodynamic state, with the pole whose real part is closest to zero governing the asymptotic behavior of the total correlation function.
@@
where $\delta_{n}$ is the argument of the associated residue $|A_{\zeta_{n}}|e^{\pm\imath\delta_{n}}$. The first term on the right-hand side of Equation (5) dominates over the second one if $\zeta_{1}<\zeta_{2}$; conversely, the second term dominates if $\zeta_{1}>\zeta_{2}$. Given a value of $\beta p$, there may exist a certain temperature at which the conditions $\zeta_{1}=\zeta_{2}$ and $\omega_{1}\neq\omega_{2}$ are satisfied. The set of such states plotted on the $T$ vs. $\beta p$ plane (or equivalently on the $T$ vs. $\rho$ plane) defines the DOC line. When this line is crossed, the wavelength of the asymptotic damped oscillations in $h(r)$ undergoes a discontinuous shift from $2\pi/\omega_{1}$ to $2\pi/\omega_{2}$ (or vice versa).
@@
For a given value of $\beta p$, there may exist a specific temperature at which the conditions $\zeta=\kappa$ and $\omega\neq 0$ are satisfied. The collection of such states, when plotted on the $T$ vs. $\beta p$ plane (or equivalently on the $T$ vs. $\rho$ plane), defines the FW line. Upon crossing this line, the nature of the decay of the total correlation function $h(r)$ transitions between damped oscillatory and monotonic behavior (or vice versa).
@@
Regardless of the sign of $\epsilon_{j}$, the leading pole at a given density, $\rho^{*}$, for the high-temperature limit ($\beta\to 0$) is given by $\zeta=\zeta_{\text{HR}}(\rho^{*})$ and $\omega=\omega_{\text{HR}}(\rho^{*})$, as shown in Appendix A.2, where the subscript HR refers to the hard rod fluid. The HR oscillation frequency satisfies $\frac{1}{2}<\omega_{\text{HR}}(\rho^{*})/2\pi<1$, with the lower and upper bounds corresponding to $\rho^{*}\to 0$ and $\rho^{*}\to 1$, respectively. On the other hand, if $\epsilon_{j}>0$ and $\rho^{*}<\lambda_{2}^{-1}$, the leading pole for the low-temperature limit ($\beta\to\infty$) is given by $\zeta=\lambda_{2}^{-1}\zeta_{\text{HR}}(\rho^{*}\lambda_{2})$ and $\omega=\lambda_{2}^{-1}\omega_{\text{HR}}(\rho^{*}\lambda_{2})$ (see Appendix A.3.1). However, the low-temperature limit for $\lambda_{2}^{-1}<\rho^{*}<1$ is more intricate, as detailed in Appendix A.3.2.